OpenStudy (anonymous):

Divide −16x^8/−8x^9 A. −2/x B. −2x C. 2x D. 2/x

OpenStudy (anonymous):

A

OpenStudy (anonymous):

\[\frac{−16x^8}{−8x^9}=\frac{−16}{−8}x^{8-9}=2x^{-1}=\frac{2}{x}\] @LaBlueEyes2013